# Multi-View Adaptive Contrastive Learning for Information Retrieval Based Fault Localization

### v2

@@ 1 Introduction
Existing approaches can be categorized into two mainstreams: 1) Spectrum-based approaches [1, 2, 3, 4], which require dynamic spectra information between different executions of the program to help locate faults. However, it is hard to collect many passing and failing execution traces in software maintenance phase. 2) Information retrieval (IR)-based approaches [5, 6], which do not require dynamic execution information, but analyze the text of bug reports and source code files to measure the relevance between them. However, there is a significant lexical gap between the natural language in bug reports and the programming language in source code files [6]. To bridge the lexical gap, researchers designed a series of semantic matching models utilizing deep neural network and IR techniques, and improved localization performance [7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
Although semantic matching based models have made some progress, the performance of IR-based approaches is often affected by the textual quality of bug reports. Specifically, it is very difficult to have satisfactory performance when insufficient textual description is provided in bug reports, even with very sophisticated models. In fact, apart from mining the relationship within an individual pair of bug report and code file, there are some other auxiliary information that can facilitate IR-based fault localization. For example, 1) the similarity relationships between bug reports. It is possible to receive duplicate or similar bug reports in a defect tracking system [17]. Previous works [5, 18, 19] show that if two bug reports are similar, they are likely to be resulted by similar buggy files. Thus, we can adopt such similarity relationships to recommend potential candidates. And 2) the co-citation relationships between code files. In many cases, multiple files are co-cited in one bug report, and are coupled in the corresponding bug fixing. Such a coupling relation may be intrinsically due to the software code dependency (for example, calling hierarchy), therefore may also exist in other bug fixing and the coupled files are located together consequently.
@@
These auxiliary information can be leveraged to better locate in the case of insufficient bug textual information [19]. However, existing studies have not deeply explored the impact of multiple implicit information on location performance, and how to optimize the adaptive extraction of valuable information from multiple sources. All the above discussion motivates us to design a new automatic fault localization approach, which utilizes auxiliary information to supplement the insufficient bug textual information, and automatically encodes different information into the representation learnings of bug reports and source code files.
Fig. 1(a) shows the report-code bug-fixing history interaction relationships modeled as a typical bipartite graph structure. In practice, there are lots of implicit relationships among bug reports and source code files, which might help better locate the bug-relevant files based on a given bug report. We integrate the implicit relationships into the bipartite graph as shown in Fig. 1(b). For example, in Fig. 1(b), as bug report $r_{1}$, the relevant buggy files are $c_{1}$,$c_{2}$,$c_{4}$, so there are explicit interactions between them (the black line). At the same time, there are co-citation relationships between $c_{1}$,$c_{2}$ and $c_{4}$ (the gray line). In addition, there exists a certain degree of similarity between $r_{1}$ and $r_{2}$ (or $r_{1}$ and $r_{3}$), as they share commonly fixed buggy files (the red line). To a certain extent, these relationships can help the localization process.
@@ 2.1 IR-based Fault Localization
IR-based approaches treat a bug report as a query and a source file as a document, and provide a recommendation list of candidate source files based on the textual similarity between the report and the source file. Early IR-based studies rely on the textual similarity between the bug reports and source code files to identify suspicious files using classical IR techniques such as Latent Dirichlet Allocation (LDA) [20, 21, 22], Vector Space Model (VSM) [5, 23], Latent Semantic Analysis (LSA) [24]. The basic intuition for adopting these text-based approaches for fault localization is that bug reports and project resources share common words, which may falter in their effectiveness when the source code contains limited common terms with a new bug report, potentially hindering their ability to accurately locate faults. Although there may not be substantial overlap in the shared common words, there is a high likelihood that they convey the same meaning semantically. Consequently, to bridge the gap between natural language and programming code, extensive research [7, 8, 9, 25, 26, 10, 11, 27, 12] has been devoted to leveraging deep learning techniques for semantic matching. Although these semantic-based approaches achieved promising performance, they directly used representation-focused models from natural language processing (NLP) tasks to learn the representation of bug reports and source files without considering their characteristics. To this end, [28, 14, 16] constructed source code as control flow graph (CFG), code knowledge graph, and code property graph (CPG) respectively, aiming to extract intricate structural information to enrich the code characteristics. Furthermore, in addition to semantic matching, several studies have integrated other techniques to enhance the performance of fault localization. For instance, Zhu et al. [29] employed a generative adversarial network (GAN) [30, 31] model to learn from historical bug-fixing records and simulate the generation of new fix proposals for localization.
@@ 2.2 Graph Neural Networks
The graph data is a type of unstructured data that consists of nodes and edges, which can accurately describe the complex interaction between nodes. In recent years, graph neural networks (GNNs) have established a deep learning framework for graph structure, which can utilize both topology structure information and node feature information. The majorities of typical GNNs, such as graph convolutional networks (GCNs) [32, 33, 34], graph attention networks (GAT) [35], graphsage [36], etc, are designed for homogeneous graphs, which disregard the variations in both the node type and edge type. In the real world, there is more than one kind of interaction between things, that is, there are multiple types of nodes and their varying edges. To model the heterogeneity of graphs, there are two types of approaches. The first type is to convert a heterogeneous graph into multiple homogeneous graphs, employ typical GNNs to aggregate messages, and then fused these information via attention mechanism [37]. The other type intends to model the heterogeneity directly, via different kinds of nodes and relations, represented by R-GCN [38]. R-GCN models the relations in graphs by employing specialized parameter matrices, which separately constructs GNNs on each relation graph. These GNN-based models have been widely used in node classification [39, 40], link prediction [41], knowledge graphs [42] and recommender systems [43, 44, 45].
For fault localization, although there are some researches [46, 28, 14, 16] have built localization models from the perspective of graph, few studies have considered the multiple interaction views between bug reports and source code files simultaneously. Inspired by the high accuracy and universality of graph neural networks in node-level, edge-level and graph-level analytical tasks [47], we develop a multi-view embedded fault localization approach based on GNNs to locate the bug-relevant files.
@@ 2.3 Contrastive Learning
Contrastive learning is a discriminant representation learning framework, which aims to learn high-quality representation via a self-supervised manner. The common motivation behind these work is the InfoMax principle [48], which we here instantiate as maximizing the mutual information (MI) between two views (i.e., report-code and report-report views, or report-code and code-code views). It learns discriminative representations through a comparative analysis of samples, distinguishing between those that are similar (positive samples) and those that are dissimilar (negative samples). This contrastive process ensures that representations of similar samples are pulled closer together in the representation space, while representations of dissimilar samples are pushed further apart, thereby enhancing the discriminative capability of the learned representations. Contrastive learning is prevalently applied to self-supervised learning tasks, enabling the training of models on unlabeled data. Remarkably, some instances of self-supervised contrastive learning models have demonstrated performance that is comparable to supervised models. Due to the good performance of contrastive learning, it has been widely used in computer vision [49, 50] and natural language processing [51, 52].
@@ 3.1 Motivating examples
Fig. 2 displays bug reports #$\#56029$ and #$\#33356$, and their relevant buggy files in ground truth.
Figure 2: Bug Reports #$\#56029$ and #$\#33356$, and their relevant source code files.
Assume that report #$\#33356$ and its ground truth are known historical information, and #$\#56029$ is a newly submitted bug report. Notably, an attempt to directly match the text information of the bug report #$\#56029$ with candidate source code files, using cosine similarity, yielded suboptimal results, with ELParser.java failing to rank among the top-10. On the one hand, the description of #$\#56029$ incorporates stack trace information rich in class names, which inadvertently introduces noise into the process of matching it with relevant source code files. On the other hand, while the description also encompasses code snippets from JSP pages, the direct textual overlap between the front-end code in JSP pages and the back-end code in ELParser.java is often limited. As a result, the combination of excessive stack trace noise and the inherent disconnect between front-end code snippets and back-end parsing logic contributed to the suboptimal ranking of ELParser.java in the context of this specific bug report. From another perspective, we can see that both reports #$\#33356$ and #$\#56029$ describe an exceptional bug related to the parsing of tag attributes within JSP pages. By exploiting the historical knowledge embodied in report #$\#33356$, which indicates that the buggy file associated with its resolution was ELParser.java, we can hypothesize that similar bugs may also involve this file [5, 18, 19]. We leverage the potential of aggregating information from similar bug reports to indirectly infer the involvement of potentially related files, thereby improving the efficiency and accuracy of locating the root cause of bugs.
Scenario 2. Fig. 3 shows six bug reports and their relevant buggy files in ground truth. Assume that the five reports on the left are historical data and #$\#68096$ is a newly submitted report. If we simply match report #$\#68096$ to each individual code file, it could be very hard to locate all the true buggy files together. However, if considering the co-citation relationships among these files in historical bug reports and build a network, we can observe the coupling connection between them. For example, according to the bug-fixing history of the five bug reports, the generated connections include $(1,2),(1,3),(2,4),(1,5),(1,6)$. Then, a co-citation graph can be constructed with the connection link between each co-cited pair of files, as shown in Fig. 3. As a consequence, we can quickly recommend these buggy files as potential candidate locations to report #$\#68096$ through the co-citation network.
@@ 3.2 Definitions
where $tex\_sim(r_{i},c_{j})=\frac{\mathbf{e}_{r_{i}}\cdot\mathbf{e}_{c_{j}}}{||\mathbf{e}_{r_{i}}||||\mathbf{e}_{c_{j}}||}$ is use to calculate a similarity score between the report $r_{i}$ and the code file $c_{j}$, which serves as the weight of the edge $<r_{i},c_{j}>$. Here, $\mathbf{e}_{r_{i}}$ and $\mathbf{e}_{c_{j}}$ represent the vector representations of $r_{i}$ and $c_{j}$ respectively, which are obtained by embedding these entities using CodeBERT [53].
@@
each value of $tex\_sim(r_{i},r_{j})=\frac{\mathbf{e}_{r_{i}}\cdot\mathbf{e}_{r_{j}}}{||\mathbf{e}_{r_{i}}||||\mathbf{e}_{r_{j}}||}$ is the weight of edge $<r_{i},r_{j}>$, which indicates the cosine similarity score of bug reports $r_{i}$ and $r_{j}$.
Definition 3 (Code-Code Co-citation Relationship). The implicit relationships between fault locations that are co-cited by the same bug reports can be further modeled as a co-citation network to help predict faults [46]. As such, we construct a code-code graph, and the adjacency matrix $\mathbf{A}_{C-C}=\{a_{cc}(c_{i},c_{j})|c\in C\}$, where
@@
each value of $co\_cited\_count(c_{i},c_{j})$ represents the number of times that the code files $c_{i}$ and $c_{j}$ were fixed by the same bug report, which serves as the weight of the edge $<c_{i},c_{j}>$.
@@ 4.3 Joint Learning Layer
The fault localization task is considered as a ranking problem and the loss function we adopt the pointwise Cross-Entropy loss [54] during the training process, which can compute a probability distribution over all candidate code files and maximize the difference between the scores of positive samples and negative samples as much as possible. For report-code interaction relationship, the observed interactions should be given a higher prediction score than those unobserved interactions. Therefore, the objective function as follows:
@@
Note that the representation space of report-code, report-report and code-code views are different. Therefore, the second task is to leverage contrastive learning to encourage graph encoders to preserve the information shared across report-code interaction view to report-report and code-code views. As shown in Fig. 4, we feed report embeddings $(\mathbf{e}_{r},\mathbf{h}_{r})$ from report-report view and report-code view into two correspond MLPs, mapping them into the same space $(\mathbf{z}^{p}_{r},\mathbf{z}^{q}_{r})$, where contrastive loss of report representations is calculated. Similarly, code embeddings $(\mathbf{e}_{c},\mathbf{h}_{c})$ from code-code view and report-code view are mapped into $(\mathbf{z}^{p}_{c},\mathbf{z}^{q}_{c})$. In order to pull the positive samples closer together in the representation space and push the negative samples further apart, for each report $r_{i}$, we treat the report-report view and report-code view of the same report as a positive pair $\{\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{i}}\}$. On the other side, we pair report $r_{i}$ with a random report $r_{j}$, and get $\{\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{j}}\}$ and $\{\mathbf{z}^{p}_{r_{j}},\mathbf{z}^{q}_{r_{i}}\}$ as negative pairs. Then the contrastive loss function promotes the consistency between the representations of the two different views, while maximizes the divergence of negative pairs. That is, $sim(\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{i}})>sim(\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{j}})$ and $sim(\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{i}})>sim(\mathbf{z}^{p}_{r_{j}},\mathbf{z}^{q}_{r_{i}})$. Analogously, the contrastive learning process for code representations follows the above steps, emphasizing consistency and divergence across code-code view and report-code view. Formally, we adopt the contrastive learning object function based on InfoNCE loss [55]:
@@
where $sim(\mathbf{z}^{p}_{i},\mathbf{z}^{q}_{j})=\frac{\mathbf{z}^{p}_{i}\cdot\mathbf{z}^{q}_{j}}{||\mathbf{z}^{p}_{i}||||\mathbf{z}^{q}_{j}||}$ measures the cosine similarity of two vectors, $|R|$ and $|C|$ are the total number of bug reports and code files. $\tau$ denotes the temperature hyper-parameter.

### v3

@@ 1 Introduction
Existing approaches can be categorized into two mainstreams: 1) Spectrum-based approaches [yoo2017human, tu2019analysis, natan2023distributed, yan2023fault], which require dynamic spectra information between different executions of the program to help locate faults. However, it is hard to collect many passing and failing execution traces in software maintenance phase. 2) Information retrieval (IR)-based approaches [zhou2012should, ye2014learning], which do not require dynamic execution information, but analyze the text of bug reports and source code files to measure the relevance between them. However, there is a significant lexical gap between the natural language in bug reports and the programming language in source code files [ye2014learning]. To bridge the lexical gap, researchers designed a series of semantic matching models utilizing deep neural network and IR techniques, and improved localization performance [ye2016word, lam2015combining, lam2017bug, gharibi2018leveraging, xiao2017improving, xiao2019improving, qi2021dreamloc, zhang2020exploiting, liang2022modeling, han2023bjxnet].
Although semantic matching based models have made some progress, the performance of IR-based approaches is often affected by the textual quality of bug reports. Specifically, it is very difficult to have satisfactory performance when insufficient textual description is provided in bug reports, even with very sophisticated models. In fact, apart from mining the relationship within an individual pair of bug report and code file, there are some other auxiliary information that can facilitate IR-based fault localization. For example, 1) the similarity relationships between bug reports. It is possible to receive duplicate or similar bug reports in a defect tracking system [chaparro2017improving]. Previous works [zhou2012should, davies2012using, li2021survey] show that if two bug reports are similar, they are likely to be resulted by similar buggy files. Thus, we can adopt such similarity relationships to recommend potential candidates. And 2) the co-citation relationships between code files. In many cases, multiple files are co-cited in one bug report, and are coupled in the corresponding bug fixing. Such a coupling relation may be intrinsically due to the software code dependency (for example, calling hierarchy), therefore may also exist in other bug fixing and the coupled files are located together consequently.
@@
These auxiliary information can be leveraged to better locate in the case of insufficient bug textual information [li2021survey]. However, existing studies have not deeply explored the impact of multiple implicit information on location performance, and how to optimize the adaptive extraction of valuable information from multiple sources. All the above discussion motivates us to design a new automatic fault localization approach, which utilizes auxiliary information to supplement the insufficient bug textual information, and automatically encodes different information into the representation learnings of bug reports and source code files.
Fig. 1(a) shows the report-code bug-fixing history interaction relationships modeled as a typical bipartite graph structure. In practice, there are lots of implicit relationships among bug reports and source code files, which might help better locate the bug-relevant files based on a given bug report. We integrate the implicit relationships into the bipartite graph as shown in Fig. 1(b). For example, in Fig. 1(b), as bug report $r_{1}$, the relevant buggy files are $c_{1}$$c_{2}$$c_{4}$, so there are explicit interactions between them (the black line). At the same time, there are co-citation relationships between $c_{1}$$c_{2}$ and $c_{4}$ (the gray line). In addition, there exists a certain degree of similarity between $r_{1}$ and $r_{2}$ (or $r_{1}$ and $r_{3}$), as they share commonly fixed buggy files (the red line). To a certain extent, these relationships can help the localization process.
@@ 2.1 IR-based Fault Localization
IR-based approaches treat a bug report as a query and a source file as a document, and provide a recommendation list of candidate source files based on the textual similarity between the report and the source file. Early IR-based studies rely on the textual similarity between the bug reports and source code files to identify suspicious files using classical IR techniques such as Latent Dirichlet Allocation (LDA) [lukins2008source, lukins2010bug, nguyen2011topic], Vector Space Model (VSM) [zhou2012should, saha2013improving], Latent Semantic Analysis (LSA) [cleary2009empirical]. The basic intuition for adopting these text-based approaches for fault localization is that bug reports and project resources share common words, which may falter in their effectiveness when the source code contains limited common terms with a new bug report, potentially hindering their ability to accurately locate faults. Although there may not be substantial overlap in the shared common words, there is a high likelihood that they convey the same meaning semantically. Consequently, to bridge the gap between natural language and programming code, extensive research [ye2016word, lam2015combining, lam2017bug, ibrahim2023explainable, van2020review, gharibi2018leveraging, xiao2017improving, church2017word2vec, xiao2019improving] has been devoted to leveraging deep learning techniques for semantic matching. Although these semantic-based approaches achieved promising performance, they directly used representation-focused models from natural language processing (NLP) tasks to learn the representation of bug reports and source files without considering their characteristics. To this end, [huo2020control, zhang2020exploiting, han2023bjxnet] constructed source code as control flow graph (CFG), code knowledge graph, and code property graph (CPG) respectively, aiming to extract intricate structural information to enrich the code characteristics. Furthermore, in addition to semantic matching, several studies have integrated other techniques to enhance the performance of fault localization. For instance, Zhu et al. [zhu2022bl] employed a generative adversarial network (GAN) [liu2019generative, creswell2018generative] model to learn from historical bug-fixing records and simulate the generation of new fix proposals for localization.
@@ 2.2 Graph Neural Networks
The graph data is a type of unstructured data that consists of nodes and edges, which can accurately describe the complex interaction between nodes. In recent years, graph neural networks (GNNs) have established a deep learning framework for graph structure, which can utilize both topology structure information and node feature information. The majorities of typical GNNs, such as graph convolutional networks (GCNs) [kipf2016semi, wu2019simplifying, chen2020simple], graph attention networks (GAT) [velivckovic2017graph], graphsage [hamilton2017inductive], etc, are designed for homogeneous graphs, which disregard the variations in both the node type and edge type. In the real world, there is more than one kind of interaction between things, that is, there are multiple types of nodes and their varying edges. To model the heterogeneity of graphs, there are two types of approaches. The first type is to convert a heterogeneous graph into multiple homogeneous graphs, employ typical GNNs to aggregate messages, and then fused these information via attention mechanism [fan2022heterogeneous]. The other type intends to model the heterogeneity directly, via different kinds of nodes and relations, represented by R-GCN [schlichtkrull2018modeling]. R-GCN models the relations in graphs by employing specialized parameter matrices, which separately constructs GNNs on each relation graph. These GNN-based models have been widely used in node classification [zeng2021gcn2defect, zhou2022software], link prediction [zhang2023iea], knowledge graphs [park2019estimating] and recommender systems [xie2021devrec, you2019attributed, song2024xgcn].
For fault localization, although there are some researches [chen2008implicit, huo2020control, zhang2020exploiting, han2023bjxnet] have built localization models from the perspective of graph, few studies have considered the multiple interaction views between bug reports and source code files simultaneously. Inspired by the high accuracy and universality of graph neural networks in node-level, edge-level and graph-level analytical tasks [wu2022graph], we develop a multi-view embedded fault localization approach based on GNNs to locate the bug-relevant files.
@@ 2.3 Contrastive Learning
Contrastive learning is a discriminant representation learning framework, which aims to learn high-quality representation via a self-supervised manner. The common motivation behind these work is the InfoMax principle [linsker1988self], which we here instantiate as maximizing the mutual information (MI) between two views (i.e., report-code and report-report views, or report-code and code-code views). It learns discriminative representations through a comparative analysis of samples, distinguishing between those that are similar (positive samples) and those that are dissimilar (negative samples). This contrastive process ensures that representations of similar samples are pulled closer together in the representation space, while representations of dissimilar samples are pushed further apart, thereby enhancing the discriminative capability of the learned representations. Contrastive learning is prevalently applied to self-supervised learning tasks, enabling the training of models on unlabeled data. Remarkably, some instances of self-supervised contrastive learning models have demonstrated performance that is comparable to supervised models. Due to the good performance of contrastive learning, it has been widely used in computer vision [tian2020contrastive, he2020momentum] and natural language processing [fu2021lrc, sun2023contrastive].
@@ 3.1 Motivating examples
Fig. 2 displays bug reports #56029\#56029 and #33356\#33356, and their relevant buggy files in ground truth.
Figure 2: Bug Reports #56029\#56029 and #33356\#33356, and their relevant source code files.
Assume that report #33356\#33356 and its ground truth are known historical information, and #56029\#56029 is a newly submitted bug report. Notably, an attempt to directly match the text information of the bug report #56029\#56029 with candidate source code files, using cosine similarity, yielded suboptimal results, with ELParser.java failing to rank among the top-10. On the one hand, the description of #56029\#56029 incorporates stack trace information rich in class names, which inadvertently introduces noise into the process of matching it with relevant source code files. On the other hand, while the description also encompasses code snippets from JSP pages, the direct textual overlap between the front-end code in JSP pages and the back-end code in ELParser.java is often limited. As a result, the combination of excessive stack trace noise and the inherent disconnect between front-end code snippets and back-end parsing logic contributed to the suboptimal ranking of ELParser.java in the context of this specific bug report. From another perspective, we can see that both reports #33356\#33356 and #56029\#56029 describe an exceptional bug related to the parsing of tag attributes within JSP pages. By exploiting the historical knowledge embodied in report #33356\#33356, which indicates that the buggy file associated with its resolution was ELParser.java, we can hypothesize that similar bugs may also involve this file [zhou2012should, davies2012using, li2021survey]. We leverage the potential of aggregating information from similar bug reports to indirectly infer the involvement of potentially related files, thereby improving the efficiency and accuracy of locating the root cause of bugs.
Scenario 2. Fig. 3 shows six bug reports and their relevant buggy files in ground truth. Assume that the five reports on the left are historical data and #68096\#68096 is a newly submitted report. If we simply match report #68096\#68096 to each individual code file, it could be very hard to locate all the true buggy files together. However, if considering the co-citation relationships among these files in historical bug reports and build a network, we can observe the coupling connection between them. For example, according to the bug-fixing history of the five bug reports, the generated connections include $(1,2),(1,3),(2,4),(1,5),(1,6)$. Then, a co-citation graph can be constructed with the connection link between each co-cited pair of files, as shown in Fig. 3. As a consequence, we can quickly recommend these buggy files as potential candidate locations to report #68096\#68096 through the co-citation network.
@@ 3.2 Definitions
[…]
@@
[…]
Definition 3 (Code-Code Co-citation Relationship). The implicit relationships between fault locations that are co-cited by the same bug reports can be further modeled as a co-citation network to help predict faults [chen2008implicit]. As such, we construct a code-code graph, and the adjacency matrix $\mathbf{A}_{C-C}=\{a_{cc}(c_{i},c_{j})|c\in C\}$, where
@@
each value of $co\_cited\_count(c_{i},c_{j})$ represents the number of times that the code files $c_{i}$ and $c_{j}$ were fixed by the same bug report, which serves as the weight of the edge $c_{i},c_{j}>$.
@@ 4.3 Joint Learning Layer
The fault localization task is considered as a ranking problem and the loss function we adopt the pointwise Cross-Entropy loss [mao2023cross] during the training process, which can compute a probability distribution over all candidate code files and maximize the difference between the scores of positive samples and negative samples as much as possible. For report-code interaction relationship, the observed interactions should be given a higher prediction score than those unobserved interactions. Therefore, the objective function as follows:
@@
Note that the representation space of report-code, report-report and code-code views are different. Therefore, the second task is to leverage contrastive learning to encourage graph encoders to preserve the information shared across report-code interaction view to report-report and code-code views. As shown in Fig. 4, we feed report embeddings $(\mathbf{e}_{r},\mathbf{h}_{r})$ from report-report view and report-code view into two correspond MLPs, mapping them into the same space $(\mathbf{z}^{p}_{r},\mathbf{z}^{q}_{r})$, where contrastive loss of report representations is calculated. Similarly, code embeddings $(\mathbf{e}_{c},\mathbf{h}_{c})$ from code-code view and report-code view are mapped into $(\mathbf{z}^{p}_{c},\mathbf{z}^{q}_{c})$. In order to pull the positive samples closer together in the representation space and push the negative samples further apart, for each report $r_{i}$, we treat the report-report view and report-code view of the same report as a positive pair $\{\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{i}}\}$. On the other side, we pair report $r_{i}$ with a random report $r_{j}$, and get $\{\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{j}}\}$ and $\{\mathbf{z}^{p}_{r_{j}},\mathbf{z}^{q}_{r_{i}}\}$ as negative pairs. Then the contrastive loss function promotes the consistency between the representations of the two different views, while maximizes the divergence of negative pairs. That is, $sim(\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{i}})>sim(\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{j}})$ and $sim(\mathbf{z}^{p}_{r_{i}},\mathbf{z}^{q}_{r_{i}})>sim(\mathbf{z}^{p}_{r_{j}},\mathbf{z}^{q}_{r_{i}})$. Analogously, the contrastive learning process for code representations follows the above steps, emphasizing consistency and divergence across code-code view and report-code view. Formally, we adopt the contrastive learning object function based on InfoNCE loss [oord2018representation]:
@@
[…]